The distance between towns A to B is equal to 48 km. The road connecting A and B is not flat and consists of segments that go either uphill or downhill only. A cyclist rode his bike from A to B and back in 3.5 hours. It is known that the cyclist was moving uphill at a speed 24 km/h. At what speed did the cyclist move downhill?
Answer by TimothyLamb(4379)

x = downhill speed
y = uphill speed = 24 kph
---
s = d/t
---
1/24 + 1/x = 3.5/48
1/x = 3.5/48 - 1/24
x = 1/(3.5/48 - 1/24)
x = 32
---
answer:
x = downhill speed = 32 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php